Jessica and Lillian had the same amount of money. jessica gave $1140.00 to charity and Lillian gave $580.00 to a different charity. In the end Lillian had 9 times as much money as Jessica. How much money did each girl have at first?
Answer by Theo(13342)

they each have x dollars.
jessica gave 1140 to charity.
lillian gave 580 to charity.

jessica was left with x - 1140 dollars.
lillian was left with x - 580 dollars.

lillian was left with 9 times the money that jessica was left with.

(x - 580) = 9 * (x - 1140)

solve for x.

simplify to get:
x - 580 = 9*x - 9*1140 which becomes:
x - 580 = 9*x - 10260

subtract x from both sides of the equation to get:

-580 = 8*x - 10260

add 10260 to both sides of the equation to get:

10260 - 580 = 8*x which becomes 9680 = 8*x

divide both sides of the equation by 8 to get:

x = 1210.

they both started with 1210.

jessica gave away 1140 to be left with 70 dollars.

lillian gave away 580 to be left with 630 dollars

lillian is left with 630 / 70 = 9 times as much as jessica.